# Bergquist and Elmachtoub

### v4

@@ 1 Introduction
In practice, however, there are many downsides to dynamic pricing. The strategy may be unappealing to customers and appear unfair (Cohen et al. 2022). In response, customers might begin to exhibit strategic behavior such as delaying their purchase, which may not be accounted for in the model. Implementing a dynamic pricing strategy also requires increased operational complexity to communicate and update prices. Moreover, a reliable demand model is needed to ensure small price changes are indeed causing intended changes in demand. To address these issues, static pricing may be an attractive alternative. Static pricing is advantageous not only because it provides transparency for the customer, it is also simpler and more tractable for service providers to implement. However, when we restrict ourselves to static pricing policies, some loss of revenue or increased congestion may have to be endured. While it is understood that static policies are optimal in the fluid limit of many stochastic models, their theoretical performance is less understood when service and arrival rates are finite. In this work, we quantify the strength of static pricing in the worst-case by furnishing universal, non-asymptotic guarantees relative to the optimal dynamic pricing policy.
@@
The static pricing policies we consider use a single price, and stop selling when the number of customers in the queue reaches a specified threshold. These threshold-static pricing strategies are simple to implement and allow us to directly trade off revenue and congestion via the threshold parameter. Note that for static policies which do not have a threshold, we show the performance can be arbitrarily poor. Our proofs are by construction, and we use the static price that results in the same expected arrival rate as the optimal policy. This construction has two key advantages: (i) the revenue rate when the state is below the threshold is always at least as good as the optimal policy due to an application of Jensen’s inequality (ii) we can directly relate our policy to the expected queue length of the optimal policy via Little’s law. However, this use of the optimal policy’s average arrival rate (denoted $\tilde{\lambda}$) is purely a proof device; in practice, one would not need to compute $\tilde{\lambda}$ or find the optimal dynamic policy. Instead, a decision-maker can search directly over the two parameters (the price and cutoff threshold) to implement a static policy.
@@ 1.1 Literature Review
    The state of the queue does not directly impact the arrival process, which we call the Low queue since it was first introduced in Low (1974a).
@@
To our knowledge, the model we consider was first introduced in Low (1974a) and followed up with Low (1974b), works in which the author studies the structure of the optimal dynamic pricing policy and gives an algorithm to compute it assuming either a finite or compact set of possible prices, respectively. Our work uses different objective functions motivated by our applications and does not require compactness of the possible prices. The Low queue, albeit restricted to a single-server, was again considered in Latouche (1980). They consider two-price policies and, for a given pair of prices, compute the optimal policy under two different criteria: either maximizing revenue rate for a given upper bound on congestion, or minimizing congestion for a given lower bound on revenue rate. This contrasts with our approach in which we optimize a linear combination of the two objectives and allow a different price for each state in the dynamic policy. More recently, Maoui et al. (2009) considers the problem of optimizing the static price in the single-server instance of our model. However, their work is focused primarily on identifying static pricing policies and studying the structure of the optimal policies; our aim in the present work is to relate the performance of static and dynamic policies. Another work which extends Low’s model is Ata and Shneorson (2006), which considers the case when the service rate is variable and can be picked by the service provider. They show numerically that, in their setting, dynamic pricing has significant gains over static pricing. Recently, Vaze and Nair (2022) considers a model closely related to the Low queue, but where the service rates (instead of the arrival rates) are taken as the decision variables.
@@
We now discuss works which consider observable queues. The foundational and seminal work in this line is Naor (1969), which introduced observable queues and analyzed them under the assumptions of deterministic customer valuations and static pricing. The most salient point emerging from their exact analysis is that the static price which maximizes social welfare is strictly smaller than the revenue-maximizing price. In Chen and Frank (2001), the authors generalize the observable model in Naor (1969) (and their work in Chen and Frank (2004)) to allow for state-dependent pricing. Their work shows that a threshold policy is optimal but does not give explicit expressions for the threshold. The work of Borgs et al. (2014) provides explicit expressions for the optimal threshold in terms of the Lambert-W function. Gilboa-Freedman and Hassin (2014) considers the gain of dynamic over static control of a queueing system under an extension of Naor’s model where some customers are assumed to cooperate with the service provider. Kim and Randhawa (2018) analyze the value of dynamic pricing in the asymptotic regime, where the arrival and service rates are taken to infinity. They show that dynamic pricing has a significant effect on reducing the impact of stochasticity on accrued revenue, though most of the benefit can be attained by simple two-price policies. Examples of related work in this line are Paschalidis and Tsitsiklis (2000), Yoon and Lewis (2004), Maglaras and Zeevi (2003), Maglaras (2006), and Cil et al. (2011).
@@
We now discuss a few works which assume the queue is unobservable. The earliest study of this case is Edelson and Hilderbrand (1975), where they show that the revenue maximizing and social welfare maximizing prices are the same. Another work in the unobservable case is Mendelson and Whang (1990) which considers a multi-class model and provides static pricing mechanisms which maximize total social welfare. Cachon and Feldman (2011) carries out an exact analysis assuming an unobservable queue with static prices to identify when a subscription-based pricing scheme (in which users pay a monthly fee for an arbitrary number of uses) may outperform a per-use approach. Again assuming static pricing, Haviv and Randhawa (2014) analyzes the loss of optimality that must be suffered when a price must be set without knowledge of the arrival rate.
@@
The preceding works and the literature more generally have mostly dealt with the case where the queue is either completely observed or totally opaque to the customer; another avenue is the question of interpolating between these two extremes by investigating optimal signaling. This question has been addressed in a few works, including Hassin and Koshman (2017) and Lingenbrink and Iyer (2019). In these works, the authors show that with respect to a certain objective, the optimal signaling policy falls in the middle of the two extremes.
@@
The foregoing works are all unified in that the underlying queueing model is essentially an $M/M/C$ queue. We continue by reviewing works which consider models slightly further afield. Besbes et al. (2022), Elmachtoub and Shi (2023) provides universal performance guarantees for static pricing policies when the model is an Erlang loss system (which has no congestion objective), and describe a real-world implementation in Besbes et al. (2020). Shakkottai et al. (2008) introduce a stylized model involving pricing and queues that is tailored for telecommunication applications and furnish a lower bound on the ratio between the revenue earned from their simple pricing rule and the maximum possible, which they denote as the “price of simplicity”. Their model assumes no disutility is incurred for long delays in service. Lee et al. (2012) shows that when such disutility is assumed to be present, the “price of simplicity” can be very high. Since in typical applications high delays are undesirable, this insight demonstrates the importance of considering congestion in these models and motivates our centering of it in the present work. Examples of other important works which analyze pricing in queueing models that are tailored for more specific applications are Gans and Savin (2007), Xu and Li (2013), Waserhole and Jost (2016), Banerjee et al. (2015), Banerjee et al. (2017), Lin et al. (2023), and Benjaafar and Shen (2023). Finally, Elmachtoub et al. (2023) show analogous bi-criteria approximations to ours but for a completely different problem regarding joint pricing and inventory control.
@@ 2 Model
We consider a model where a service provider seeks to optimize a linear combination of the expected revenue and the expected congestion of the system. There are $C$ servers that serve price-sensitive customers in a first-come, first-serve manner. Customers arrive according to a Poisson process with rate $\Lambda>0$. Each customer has a valuation drawn i.i.d. from a distribution $F$. We make the standard assumption in the revenue management literature that $F$ is a regular distribution (also known as Myerson’s regularity, see Myerson (1981)). When a customer arrives, the provider offers service at some price $p$ which may depend on the state of the system; the customer decides to join the queue if their valuation is at least $p$. This gives rise to the effective arrival rate of customers, which we denote by $\lambda(p):=\Lambda(1-F(p))$. We assume that there is a one-to-one correspondence between prices $p$ and effective arrival rates $\lambda$ so that $\lambda(p)$ has a unique inverse, denoted by $p(\lambda)$. Thus we can equivalently view the effective arrival rates $\lambda$ as the decision variables, which simplifies our analysis. By our assumption that $F$ is regular, we have that $\lambda p(\lambda)$ is concave (Ewerhart 2013).
@@
[…]
The set of admissible policies $\mathbf{\Pi}$ is the set of non-anticipating policies. An admissible policy $\pi$ may be represented by a vector of arrival rates $\{\lambda_{0},\lambda_{1},\lambda_{2},\dots\}$. From the memoryless property, it is sufficient to let the number of customers in the system represent the state of the system. When there are $i$ customers in the system, the price is set to $p(\lambda_{i})$. In some instances, specifically when $\Lambda\geq 1$, some admissible policies induce unstable queues. To be precise, if we write $\mu_{i}=\min(i,C)$, and define $b_{n}=\Pi_{i=1}^{n}(\lambda_{i-1}/\mu_{i})$, then the resulting queue is stable if and only if $\sum_{n=0}^{\infty}b_{n}$ converges (see e.g. pg. 254 of Wolff (1989)). Otherwise, the queue increases without bound. However, we can restrict ourselves, without loss of optimality, to considering policies which induce stable queues, as those which do not induce stability incur infinitely large costs. The stationary probabilities of a policy $\pi\in\mathbf{\Pi}$ are denoted by $\mathbb{P}_{i}(\pi)$ for each state $i\in\{0,1,2\dots\}$. We denote the optimal policy by $\pi^{*}$ with corresponding optimal arrival rates of $\lambda_{0}^{*},\lambda_{1}^{*},\lambda_{2}^{*},\ldots$ and steady-state probabilities $\mathbb{P}_{0}(\pi^{*}),\mathbb{P}_{1}(\pi^{*}),\mathbb{P}_{2}(\pi^{*}),\ldots$.
The static policies we consider are those non-anticipating policies which fix a price up to a certain occupancy threshold (potentially infinite) and then disallow arrivals after that point. In other words, we consider the class of policies that fix an arrival rate $\lambda$ in every state at or below a certain cutoff point $\gamma\in\{0,1,2,\dots\}$ and blocks arrivals in state $\gamma+1$ and higher by setting the arrival rates to 00 (infinite price). An arbitrary static policy is notated $\pi^{\lambda,\gamma}$, which refers to the policy which picks the arrival rate $\lambda$ for customers whenever there are at most $\gamma$ customers in the system, and otherwise sets all arrival rates to 00.
@@ 2.1 Objectives
where $\gamma=\tfrac{1}{1+\Lambda}$ dominates the total transition rate out of state $i$ (to states $i-1,i,i+1$), and $\eta$ is the optimal average profit.
@@ 2.2 Little’s Law and Static Policies
Our key proof idea is to use this average arrival rate to construct good static policies. In any state in which our static policy sells, we pick the arrival rate $\tilde{\lambda}$. We note that this differs from the static policy used in Besbes et al. (2022) to attain performance guarantees for the Erlang loss model. Their static policy was also derived from the optimal policy although it was based on a modification of the expected arrival rate, and the analysis is quite different than ours due to the mixed objective function $\mathcal{Z}(\pi)$ we study.
@@ 3 Single-server Static Pricing Guarantees
The fact that we can provide exact [^1] The exact, closed-form expression for $g(2)$ involves the root of a quartic and would take a few pages to relay. values for the function $g(\gamma)$ is a stroke of some good luck. As will be seen in the proof, finding these worst-case cost bounds involves solving an optimization problem over $\tilde{\lambda}\in[0,1]$. For $30<\gamma<$ 3, the maximizer lies strictly within the interval $[0,1]$; for $\gamma\geq 3$, the function is monotonically increasing on the interval and hence the maximum occurs when $\tilde{\lambda}=1$. This is fortunate because finding the unconstrained maxima of the relevant ratio involves finding the roots higher and higher degree polynomials: for $\gamma=1$, a degree $2$ polynomial arises, for $\gamma=2$, a degree $4$ polynomial arises, and for general $\gamma$ a degree $2\gamma$ polynomial arises. If the unconstrained maximum when $\gamma=3$ were to lie in the interval $[0,1]$, it is unlikely we could find a closed-form expression.
@@
where $\prod_{k=0}^{-1}\lambda_{k}=1$. For policies which induce a null recurrent or transient process, these probabilities are not defined. In fact, by applying Theorem 2 of Karlin and McGregor (1957), we get that the induced process is ergodic precisely when the denominator of Equation (4) is finite. Since any policies which do not induce ergodic processes incur arbitrarily large costs, we assume our policies are such that $\sum_{n=0}^{\infty}\prod_{k=0}^{n}\lambda_{k}<\infty$.
@@
We make one more useful observation. Consider the static policy $\pi^{\lambda,\gamma}$ which sets an arrival rate $\lambda$ for all states below and including $\gamma$, and cuts off arrivals by setting all others to 00. When we fix such a policy, the queueing part of the model becomes a classical $M/M/1/\gamma+1$ model with service rate $1$ and arrival rate $\lambda$. Thus we get the following well-known form of the stationary probabilities (see e.g. Allen (1990)) which will be used in our arguments

### v5

@@ 1 Introduction
In practice, however, there are many downsides to dynamic pricing. The strategy may be unappealing to customers and appear unfair (cohen2022price). In response, customers might begin to exhibit strategic behavior such as delaying their purchase, which may not be accounted for in the model. Implementing a dynamic pricing strategy also requires increased operational complexity to communicate and update prices. Moreover, a reliable demand model is needed to ensure small price changes are indeed causing intended changes in demand. To address these issues, static pricing may be an attractive alternative. Static pricing is advantageous not only because it provides transparency for the customer, it is also simpler and more tractable for service providers to implement. However, when we restrict ourselves to static pricing policies, some loss of revenue or increased congestion may have to be endured. While it is understood that static policies are optimal in the fluid limit of many stochastic models, their theoretical performance is less understood when service and arrival rates are finite. In this work, we quantify the strength of static pricing in the worst-case by furnishing universal, non-asymptotic guarantees relative to the optimal dynamic pricing policy.
@@
The static pricing policies we consider use a single price, and stop selling when the number of customers in the queue reaches a specified threshold. These threshold-static pricing strategies are simple to implement and allow us to directly trade off revenue and congestion via the threshold parameter. Note that for static policies which do not have a threshold, we show the performance can be arbitrarily poor. We emphasize that our guarantees are universal, meaning that they do not depend on the instance and are not asymptotic in nature. Our proofs are by construction, and we use the static price that results in the same expected arrival rate as the optimal policy. This construction has two key advantages: (i) the revenue rate when the state is below the threshold is always at least as good as the optimal policy due to an application of Jensen’s inequality (ii) we can directly relate our policy to the expected queue length of the optimal policy via Little’s law. However, this use of the optimal policy’s average arrival rate (denoted $\tilde{\lambda}$) is purely a proof device; in practice, one would not need to compute $\tilde{\lambda}$ or find the optimal dynamic policy. Instead, a decision-maker can search directly over the two parameters (the price and cutoff threshold) to implement a static policy.
@@ 1.1 Literature Review
    The state of the queue does not directly impact the arrival process, which we call the Low queue since it was first introduced in 10.2307/169504.
@@
To our knowledge, the model we consider was first introduced in 10.2307/169504 and followed up with 5391267, works in which the author studies the structure of the optimal dynamic pricing policy and gives an algorithm to compute it assuming either a finite or compact set of possible prices, respectively. Our work uses different objective functions motivated by our applications and does not require compactness of the possible prices. The Low queue, albeit restricted to a single-server, was again considered in LATOUCHE1980203. They consider two-price policies and, for a given pair of prices, compute the optimal policy under two different criteria: either maximizing revenue rate for a given upper bound on congestion, or minimizing congestion for a given lower bound on revenue rate. This contrasts with our approach in which we optimize a linear combination of the two objectives and allow a different price for each state in the dynamic policy. More recently, MAOUI2009912 considers the problem of optimizing the static price in the single-server instance of our model. However, their work is focused primarily on identifying static pricing policies and studying the structure of the optimal policies; our aim in the present work is to relate the performance of static and dynamic policies. Another work which extends Low’s model is doi:10.1287/mnsc.1060.0587, which considers the case when the service rate is variable and can be picked by the service provider. They show numerically that, in their setting, dynamic pricing has significant gains over static pricing. Recently, vaze2022nonasymptotic considers a model closely related to the Low queue, but where the service rates (instead of the arrival rates) are taken as the decision variables.
@@
We now discuss works which consider observable queues. The foundational and seminal work in this line is 10.2307/1909200, which introduced observable queues and analyzed them under the assumptions of deterministic customer valuations and static pricing. The most salient point emerging from their exact analysis is that the static price which maximizes social welfare is strictly smaller than the revenue-maximizing price. In doi:10.1080/07408170108936878, the authors generalize the observable model in 10.2307/1909200 (and their work in articleChenFrank2) to allow for state-dependent pricing. Their work shows that a threshold policy is optimal but does not give explicit expressions for the threshold. The work of borgs_chayes_doroudi_harchol-balter_xu_2014 provides explicit expressions for the optimal threshold in terms of the Lambert-W function. GILBOAFREEDMAN2014217 considers the gain of dynamic over static control of a queueing system under an extension of Naor’s model where some customers are assumed to cooperate with the service provider. doi:10.1287/opre.2017.1668 analyze the value of dynamic pricing in the asymptotic regime, where the arrival and service rates are taken to infinity. They show that dynamic pricing has a significant effect on reducing the impact of stochasticity on accrued revenue, though most of the benefit can be attained by simple two-price policies. Examples of related work in this line are 842140, articleyoon, doi:10.1287/mnsc.49.8.1018.16402, doi:10.1287/opre.1060.0305, and articleCil.
@@
We now discuss a few works which assume the queue is unobservable. The earliest study of this case is 10.2307/1913415, where they show that the revenue maximizing and social welfare maximizing prices are the same. Another work in the unobservable case is 10.2307/171046 which considers a multi-class model and provides static pricing mechanisms which maximize total social welfare. RePEc:inm:ormsom:v:13:y:2011:i:2:p:244-260 carries out an exact analysis assuming an unobservable queue with static prices to identify when a subscription-based pricing scheme (in which users pay a monthly fee for an arbitrary number of uses) may outperform a per-use approach. Again assuming static pricing, doi:10.1287/msom.2014.0479 analyzes the loss of optimality that must be suffered when a price must be set without knowledge of the arrival rate.
@@
The preceding works and the literature more generally have mostly dealt with the case where the queue is either completely observed or totally opaque to the customer; another avenue is the question of interpolating between these two extremes by investigating optimal signaling. This question has been addressed in a few works, including HASSIN2017436 and doi:10.1287/opre.2018.1819. In these works, the authors show that with respect to a certain objective, the optimal signaling policy falls in the middle of the two extremes.
@@
The foregoing works are all unified in that the underlying queueing model is essentially an $M/M/C$ queue. We continue by reviewing works which consider models slightly further afield. doi:10.1287/opre.2020.2054, elmachtoub2023power provides universal performance guarantees for static pricing policies when the model is an Erlang loss system (which has no congestion objective), and describe a real-world implementation in besbes2020pricing. 4604751 introduce a stylized model involving pricing and queues that is tailored for telecommunication applications and furnish a lower bound on the ratio between the revenue earned from their simple pricing rule and the maximum possible, which they denote as the “price of simplicity”. Their model assumes no disutility is incurred for long delays in service. 6354274 shows that when such disutility is assumed to be present, the “price of simplicity” can be very high. Since in typical applications high delays are undesirable, this insight demonstrates the importance of considering congestion in these models and motivates our centering of it in the present work. Examples of other important works which analyze pricing in queueing models that are tailored for more specific applications are doi:10.1287/mnsc.1060.0651, 6671562, articleWasherholeJost, 10.1145/2764468.2764527, 10.1145/3033274.3085099, doi:10.1287/mnsc.2022.4474, and benjaafar2023pricing. Finally, elmachtoub2023simple show analogous bi-criteria approximations to ours but for a completely different problem regarding joint pricing and inventory control.
@@ 2 Model
We consider a model where a service provider seeks to optimize a linear combination of the expected revenue and the expected congestion of the system. There are $C$ servers that serve price-sensitive customers in a first-come, first-serve manner. Customers arrive according to a Poisson process with rate $\Lambda>0$. Each customer has a valuation drawn i.i.d. from a distribution $F$. We make the standard assumption in the revenue management literature that $F$ is a regular distribution (also known as Myerson’s regularity, see doi:10.1287/moor.6.1.58). When a customer arrives, the provider offers service at some price $p$ which may depend on the state of the system; the customer decides to join the queue if their valuation is at least $p$. This gives rise to the effective arrival rate of customers, which we denote by $\lambda(p):=\Lambda(1-F(p))$. We assume that there is a one-to-one correspondence between prices $p$ and effective arrival rates $\lambda$ so that $\lambda(p)$ has a unique inverse, denoted by $p(\lambda)$. Thus we can equivalently view the effective arrival rates $\lambda$ as the decision variables, which simplifies our analysis. By our assumption that $F$ is regular, we have that $\lambda p(\lambda)$ is concave (Ewerhart2013RegularTD).
@@
The parameters for our model depend upon the units we choose for currency and time. By picking the units in a certain way, we can transform an arbitrary instance it into one with $\mu=c=1$ without impacting the approximation ratios of interest. Specifically, given an arbitrary instance, we can translate into an instance with $\mu=c=1$ by measuring all monetary values in units of $c$ and time in units of $\mu$. For example, if $\Lambda=2\text{ per hour}$, $\mu=\text{ 1/2 per hour}$, $c=\mathdollar 50\text{ per hour}$, and $\lambda(p)=2-p/100\text{ per hour}$ (where $p$ is in dollars), we can instead consider the instance with $\Lambda^{\prime}=4\text{ per half hour}$, $\mu^{\prime}=1\text{ per half hour}$, $c^{\prime}=1\text{ hundred dollars per half hour}$, and $\lambda^{\prime}(p)=4-2p\text{ per half hour}$, (where $p$ is now measured in hundreds of dollars). In general, the “nominal” revenue, cost, and profit rates garnered by a policy change by a factor of $\frac{1}{\mu c}$ when making this translation. In the performance ratios that the present work is concerned with, this factor is present in both the numerator and denominator and thus cancels out. Hence, we assume without loss of generality that $\mu=c=1$ in all proofs.
The set of admissible policies $\mathbf{\Pi}$ is the set of non-anticipating policies. An admissible policy $\pi$ may be represented by a vector of arrival rates $\{\lambda_{0},\lambda_{1},\lambda_{2},\dots\}$. From the memoryless property, it is sufficient to let the number of customers in the system represent the state of the system. When there are $i$ customers in the system, the price is set to $p(\lambda_{i})$. In some instances, specifically when $\Lambda\geq 1$, some admissible policies induce unstable queues. To be precise, if we write $\mu_{i}=\min(i,C)$, and define $b_{n}=\Pi_{i=1}^{n}(\lambda_{i-1}/\mu_{i})$, then the resulting queue is stable if and only if $\sum_{n=0}^{\infty}b_{n}$ converges (see e.g. pg. 254 of wolff1989stochastic). Otherwise, the queue increases without bound. However, we can restrict ourselves, without loss of optimality, to considering policies which induce stable queues, as those which do not induce stability incur infinitely large costs. The stationary probabilities of a policy $\pi\in\mathbf{\Pi}$ are denoted by $\mathbb{P}_{i}(\pi)$ for each state $i\in\{0,1,2\dots\}$. We denote the optimal policy by $\pi^{*}$ with corresponding optimal arrival rates of $\lambda_{0}^{*},\lambda_{1}^{*},\lambda_{2}^{*},\ldots$ and steady-state probabilities $\mathbb{P}_{0}(\pi^{*}),\mathbb{P}_{1}(\pi^{*}),\mathbb{P}_{2}(\pi^{*}),\ldots$.
The static policies we consider are those non-anticipating policies which fix a price up to a certain occupancy threshold (potentially infinite) and then disallow arrivals after that point. In other words, we consider the class of policies that fix an arrival rate $\lambda$ in every state at or below a certain cutoff point $\gamma\in\{0,1,2,\dots\}$ and blocks arrivals in state $\gamma+1$ and higher by setting the arrival rates to 0 (infinite price). An arbitrary static policy is notated $\pi^{\lambda,\gamma}$, which refers to the policy which picks the arrival rate $\lambda$ for customers whenever there are at most $\gamma$ customers in the system, and otherwise sets all arrival rates to 0.
@@ 2.1 Objectives
where $\gamma=\tfrac{1}{1+\Lambda}$ dominates the total transition rate out of state $i$ (to states i1,i,i+1i-1,i,i+1), and $\eta$ is the optimal average profit.
@@ 2.2 Little’s Law and Static Policies
Our key proof idea is to use this average arrival rate to construct good static policies. In any state in which our static policy sells, we pick the arrival rate $\tilde{\lambda}$. We note that this differs from the static policy used in doi:10.1287/opre.2020.2054 to attain performance guarantees for the Erlang loss model. Their static policy was also derived from the optimal policy although it was based on a modification of the expected arrival rate, and the analysis is quite different than ours due to the mixed objective function $\mathcal{Z}(\pi)$ we study.
@@ 3 Single-server Static Pricing Guarantees
The fact that we can provide exact [^1] The exact, closed-form expression for $g(2)$ involves the root of a quartic and would take a few pages to relay. values for the function $g(\gamma)$ is a stroke of some good luck. As will be seen in the proof, finding these worst-case cost bounds involves solving an optimization problem over $\tilde{\lambda}\in[0,1]$. For $0<\gamma<3$, the maximizer lies strictly within the interval $[0,1]$; for $\gamma\geq 3$, the function is monotonically increasing on the interval and hence the maximum occurs when $\tilde{\lambda}=1$. This is fortunate because finding the unconstrained maxima of the relevant ratio involves finding the roots higher and higher degree polynomials: for $\gamma=1$, a degree $2$ polynomial arises, for $\gamma=2$, a degree $4$ polynomial arises, and for general $\gamma$ a degree $2\gamma$ polynomial arises. If the unconstrained maximum when $\gamma=3$ were to lie in the interval $[0,1]$, it is unlikely we could find a closed-form expression.
@@
where $\prod_{k=0}^{-1}\lambda_{k}=1$. For policies which induce a null recurrent or transient process, these probabilities are not defined. In fact, by applying Theorem 2 of Karlin1957TheCO, we get that the induced process is ergodic precisely when the denominator of Equation (4) is finite. Since any policies which do not induce ergodic processes incur arbitrarily large costs, we assume our policies are such that $\sum_{n=0}^{\infty}\prod_{k=0}^{n}\lambda_{k}<\infty$.
@@
We make one more useful observation. Consider the static policy $\pi^{\lambda,\gamma}$ which sets an arrival rate $\lambda$ for all states below and including $\gamma$, and cuts off arrivals by setting all others to 0. When we fix such a policy, the queueing part of the model becomes a classical $M/M/1/\gamma+1$ model with service rate $1$ and arrival rate $\lambda$. Thus we get the following well-known form of the stationary probabilities (see e.g. 1990665) which will be used in our arguments